OpenStudy (javyb13):

the equation of a parabola is 12y=(x-1)^2-48. Identify the vertex focus and the directrix of the parabola

OpenStudy (legomyego180):

You still there?

OpenStudy (javyb13):

yes

ILovePuppiesLol (ilovepuppieslol):

@legomyego180

OpenStudy (javyb13):

this is the last question on my quiz i really need help

OpenStudy (legomyego180):

Are you familiar with vertex form?

OpenStudy (javyb13):

uhh not really this is really confusing to me

OpenStudy (legomyego180):

It's ok. To find your vertex when the equation is written like this look at whatever x is being subtracted by. Can you tell what it is in this case?

OpenStudy (javyb13):

1?

OpenStudy (legomyego180):

Keep in mind this only works when the equation is in this form: \[y=(x-a)^2-c\] where a and c are any number. Yes, 1. This means that the x-value of your vertex is one, so we have figured out half of it. (1,y) Still need to find the y value. Any ideas on how to do that?

OpenStudy (javyb13):

divide 12 y from each side?

OpenStudy (legomyego180):

Sure, you can do that. But if you have an x value, and you want to find its corresponding y vaule, you can just plug it back into the function. So plug x=1 into your function and see what you get for y. That will give you both points of your vertex. Does that make sense?

OpenStudy (javyb13):

yes ok let me try that

OpenStudy (javyb13):

i got 4

OpenStudy (legomyego180):

-4

OpenStudy (javyb13):

oh ok i probably missed a step or something

OpenStudy (javyb13):

so thats the vertex?

OpenStudy (legomyego180):

(1,-4) = vertex. Keep in mind if this was something like (x+1) instead of (x-1) your x value for the vertex would be -1. Its whatever makes the thing inside the parentheeses zero.

OpenStudy (javyb13):

ok got it

OpenStudy (javyb13):

what about the rest

OpenStudy (legomyego180):

One sec. I havent done this in a while, so I'm learning this as I teach it

OpenStudy (javyb13):

ok

OpenStudy (legomyego180):

https://www.youtube.com/watch?v=04nBaKx9wiM

OpenStudy (legomyego180):

I reccomend watching that all the way through.

OpenStudy (javyb13):

ok thanks

OpenStudy (legomyego180):

So first thing lets rewrite \[12y=(x-1)^2-48\] as \[y+48=\frac{ 1 }{ 12 }(x-1)^2\] because w want to get the x^2 value by itself.

OpenStudy (legomyego180):

Actually, that should just be \[12y+48=(x-1)^2\]

OpenStudy (legomyego180):

Then: 4P=12 P=3 So add 3 to your vertex's y value to get your Focus, and subtract P to get your directrix. Sorry, im kind of lousy at this stuff

OpenStudy (legomyego180):

But it is the correct answer.

OpenStudy (javyb13):

thanks its ok!